OpenStudy (anonymous):

cos x + 2 cos x sin x = 0

OpenStudy (anonymous):

factor out a cos, so cos(2sin+1)

OpenStudy (anonymous):

then split it up, ya get cos x=0 and 2sin x+1=0, two different equations

OpenStudy (anonymous):

then what?

OpenStudy (jhannybean):

so 1. cos x = 0 therefore \(x = \cos^{-1} (0) =?\) 2. 2sin(x) +1 = 0 therefore \(\sin(x) = -\frac{1}{2} \implies x = \sin^{-1} \left(\frac{1}{2}\right) = ?\)

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

np